OpenStudy (anonymous):

slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = x + 2.

OpenStudy (anonymous):

A clue: if it is perpendicular it is going to be of the form y = -x + c Can you finish it with what you have?

OpenStudy (anonymous):

yes most likely.. thank you dmot7291

OpenStudy (anonymous):

What i would do is find the line itself by using the point-slope formula. That is: (y-b)=m(x-a) where a is -5 and b is-1 and m is -1 since you are looking for the perpendicular line. Then you would solve for the variable y to get y=-x+5-1 which simplifies to y=-x+4. Make sense?

OpenStudy (anonymous):

It makes sense, but I think you have gone wrong with a minus sign somewhere! y-b = m(x-a) y -(-1) = -1(x - (-5)) y +1 = - (x +5) y +1 = -x-5 y = -x -6

OpenStudy (anonymous):

hmm probably but the fact that you were able to point that out means you know what you're doing now so good job!!

OpenStudy (anonymous):

:) Just be careful with those minus signs, it is how they try to trick you!